The 95% confidence interval for the variance and standard deviation for the sugar content in an ice cream (in mg) if the sample of nine servings has a variance of 36.


1
Expert's answer
2021-05-07T12:28:13-0400

Consider that the sample mean "= 10 = \bar x"

Sample variance "= s = 36"

Sample size = "9"

z score for 95% confidence "= 1.96"

Hence,

"CI = \bar x \pm z \dfrac{s}{\sqrt{n}}"


"CI = 10 \pm 1.96 \times \dfrac {36}{3}"


"CI = 10 \pm 23.52"